# Note on two-point mean square displacement

### v1

@@ V.1 Relaxation function
Without loss of generality, we assume $n_{1}>n_{2}$ and set ${\vec{r}}(n_{1},t_{1})=(d_{0}/\sqrt{3})(1,1,1)$ and ${\vec{r}}(n_{2},t_{1})=(0,0,0)$. This implies the “initial condition” at time $t$ for each Cartesian component ($\alpha=x,y,z$)
$\langle r_{\alpha}(n,t_{1})\rangle_{d_{0}}$ $=$ $\left\{\begin{array}[]{ll}\frac{d_{0}}{3}&(n_{1}<n)\\ \frac{n-n_{2}}{n_{1}-n_{2}}\frac{d_{0}}{3}&(n_{2}<n\leq n_{1})\\ 0&(n\leq n_{2})\\ \end{array}\right.$ (63)
@@
[…]
@@
$\langle r_{\alpha}(n_{1},t_{1}+\tau)\rangle_{d_{0}}=$ $\frac{d_{0}}{2\sqrt{3}}\left[\mathrm{erf}{\left(\sqrt{\frac{\tau_{m}}{4\tau}}\right)}+1\right]$ (65)
@@
$\langle r_{\alpha}(n_{2},t_{1}+\tau)\rangle_{d_{0}}=$ $\frac{d_{0}}{2\sqrt{3}}\left[1-\mathrm{erf}{\left(\sqrt{\frac{\tau_{m}}{4\tau}}\right)}\right]$ (66)
@@
where $\tau_{m}=\tau_{0}m^{2}$ with $m=n_{1}-n_{2}$. Equations (65)and (66) together with the definition of relaxation function in Eq. (28) leads to Eq. (30).
V.2 Simulation model

### v2

@@ VI.1 Relaxation function
Without loss of generality, we assume $n_{1}>n_{2}$ and set ${\vec{r}}(n_{1},t_{1})=(d_{0}/\sqrt{3})(1,1,1)$ and ${\vec{r}}(n_{2},t_{1})=(0,0,0)$. This implies the “initial condition” at time $t_{1}$ for each Cartesian component ($\alpha=x,y,z$)
[…]
@@
$\langle r_{\alpha}(n,t_{1}+\tau)\rangle_{d_{0}}=\int dn^{\prime}\ G(n-n^{\prime},\tau)\langle r_{\alpha}(n^{\prime},t_{1})\rangle_{d_{0}}$ (68)
@@
$\langle r_{\alpha}(n_{1},t_{1}+\tau)\rangle_{d_{0}}=$ $\frac{d_{0}}{2\sqrt{3}}\left[\mathrm{erf}{\left(\sqrt{\frac{\tau_{m}}{4\tau}}\right)}+1\right]$
@@
$\langle r_{\alpha}(n_{2},t_{1}+\tau)\rangle_{d_{0}}=$ $\frac{d_{0}}{2\sqrt{3}}\left[1-\mathrm{erf}{\left(\sqrt{\frac{\tau_{m}}{4\tau}}\right)}\right]$
@@
where $\tau_{m}=\tau_{0}m^{2}$ with $m=n_{1}-n_{2}$. Equations (LABEL:r_1_ave)and (LABEL:r_2_ave) together with the definition of relaxation function in Eq. (31) leads to Eq. (33).
VI.2 Simulation of probe dynamics in moving substrate
We consider the motion of probe particle undergoing fractional Brownian motion (fBm). Denoting the probe trajectory as ${\vec{x}}(t)$, fBm is characterized by the stationary increment $\Delta{\vec{x}}(\tau)={\vec{x}}(t_{0}+\tau)-{\vec{x}}(t_{0})$, whose auto-correlation is given by
$\langle\Delta x(\tau)\Delta x(\tau^{\prime})\rangle=K_{\alpha}(\tau^{\alpha}+\tau^{\prime\alpha}-|\tau-\tau^{\prime}|^{\alpha})$ (71)
for each Cartesian component with $\alpha$ being the twice of the so-called Hurst exponent, the time and the length to be measured in appropriate units. If we assume fBm as a model of a tagged monomer in polymer (see Sec. IV), we may identify monomeric time and length scales ($\tau_{0}$ and $a$, respectively) as these units.
To describe the translational motion of the substrate, let ${\vec{X}}(t)$ be the position of its center of mass, and assume a diffusive dynamics for each component
$\frac{dX(t)}{dt}=\sqrt{2D_{T}}\ \xi_{T}(t)$ (72)
with $\langle\xi_{T}(t)\rangle=0$, $\langle\xi_{T}(t)\xi_{T}(t^{\prime})\rangle=\delta(t-t^{\prime})$ and the translational diffusion coefficient $D_{T}$.
Assuming the substrate to be isotropic, its rotational motion is described by a single angular velocity vector ${\vec{\Omega}}(t)$. Similarly to the translational motion, we assume a diffusive dynamics for each component
$\Omega(t)=\sqrt{2D_{R}}\ \xi_{R}(t)$ (73)
with $\langle\xi_{R}(t)\rangle=0$, $\langle\xi_{R}(t)\xi_{R}(t^{\prime})\rangle=\delta(t-t^{\prime})$ and the rotational diffusion coefficient $D_{R}$.
To set a scale, we introduce a characteristic length $R_{0}$, which corresponds to the spatial size of the substrate. At the start of the simulation, we set the center of mass of the substrate to be the origin and place a pair of probes at random positions within a sphere with radius $cR_{0}$ centred at origin, where $c\simeq 1$ is a constant. At each time step, each probe initially performs its own fBm displacement, then is followed by the extrinsic displacements due to substrate translation and rotation. We employ Rodrigues’ rotation formula to compute the rotational displacement from ${\vec{\Omega}}(t)$.
Our purpose in this paper is to see the regime, where the confinement effect is not relevant. In Fig. 3, we set $\alpha=0.5,K_{\alpha}=0.5,D_{T}=0.1,D_{R}=10^{-3}$, and $R_{0}=20$, $c=0.5$. Then, on the time scale shown in Fig. 3 (at least, up to $\tau=10^{2}$), the typical displacement of each probe due to its intrinsic motion (calculated as the square root of MSD) is less than $R_{0}$, thus ensuring our aim. Note that if the viscous resistance to the probe particle and that to the substrate were characterized by the same viscosity, the translational and rotational diffusion coefficients of the substrate are, respectively, $D_{T}=a/R_{0}$ and $D_{R}=(a/R_{0})^{3}$, where Eqs (72) and (73) are made dimensionless with the unit length $a$ and time $\tau_{0}$. However, this is not likely the case, for instance, in the chromatin experiment, where the viscosity would be scale-dependent as is usually expected for systems with mesoscopic structures.
As shown in Fig. 3 in main text, the observed ${\rm MSD}^{o}(\tau)$ (and also ${\rm MSCV}^{(o)}(\tau)$ on rotating substrate) fails to capture the intrinsic probe mobioity. For sub-diffusional probe ($\alpha<1$), the contribution of the substrate motion is negligible on short time scale, but it eventually dominates on longer time scale. The cross over time scale $\tau_{c}$ can be estimated by comparing the intrinsic probe displacement $\sim\tau^{\alpha}$ with the extrinsic substrate contribution $\sim D_{T}\tau$ or $\sim R_{0}^{2}D_{R}\tau$ for translational or ratational contribution. One thus finds $\tau_{c}=\max\{\tau_{c,T},\tau_{c,R}\}$ with $\tau_{c,T}\sim D_{T}^{1/(\alpha-1)}$, $\tau_{c,R}\sim(R_{0}^{2}D_{R})^{1/(\alpha-1)}$. For probes with $\alpha=1$, such a crossover is not expected, where the motion of the substrate leads to a higher apparent diffusion coefficient.
VI.3 Model of polymer simulation